Prove that a countable product of connected space is connected
1
Expert's answer
2021-08-05T08:16:36-0400

Solution

Let A and B be connected spaces such that a map

F:A×B"\to" {0,1}

To prove that A×B is connected we have to show that F is a constant.

Let

"(a_1,b_1)\in A×B"

Now ,we define

g:A"\to" {0,1}

g:a"\to" f(a,b1)

and

h:B"\to" {0,1}

h:b"\to" f(a1,b)

Since A and B are connected, h and g must be constant, hence

f(a,b1)=f(a1,b)=f(a1,b1) for all (a,b)"\in" A×B

If we fix some other;

"(a_2,b_2)\in A×B"

But "(a_1,b_1)\ne (a_2,b_2)"

We obtain

f(a,b2)=f(a2,b)=f(a2,b2) for all (a,b)"\in" A×B

In

f(a1,b)=f(a1,b1), take b=b2 ,we have

f(a1,b1)=f(a1,b2)

In

f(a,b2)=f(a2,b2), take a=a1, we have

f(a1,b2)=f(a2,b2)

Hence

f(a1,b1)=f(a2,b2)

Since

(a1,b1) and (a2,b2) are arbitrary;

f(a1,b1)=f(a2,b2) for all (a1,b1),(a2,b2) "\in" A×B

Therefore, F is a constant function.

Hence A×B is connected.